# GPU-based compressible lattice Boltzmann simulations on non-uniform grids using standard C++ parallelism: From best practices to aerodynamics, aeroacoustics and supersonic flow simulations

### v2

@@ 6.3 Viscous flow past a NACA0012 airfoil at transonic speeds
To tackle this configuration, we use a simulation domain of size $\left[-15C,15C\right]^{2}$with $C$ the airfoil chord- that is bounded by three equilibrium Dirichlet boundary conditions (left, top, and bottom) and one Neumann condition (right). As a best practice from previous works [67, 51], we start with a uniform flow at a freestream Mach number lower than the targeted one. The freestream Mach number imposed by the Dirichlet boundary conditions is then rapidly increased (over one convective time) to its nominal value. This notably allows to avoid convergence issues that were encountered with the Newton-Raphson algorithm at the beginning of the simulation for the D2Q13-LBM. When it comes to the mesh generation, four rectangular and three NACA-shaped refinement patches surround the airfoil. The finest refinement patch aims at (1) better capturing the boundary layer development at the leading edge of the airfoil, and (2) reducing pressure oscillations at the wall that are partly induced by the staircase nature of the bounce-back approach. Hereafter, simulation data are outputted at $t=5t_{c}^{\mathrm{domain}}$, so that the transient flow is completely evacuated from the simulation domain, and with the characteristic time being $t_{c}^{\mathrm{domain}}=30C/(\mathrm{Ma}\sqrt{\gamma_{r}T_{0}})$. If one is only interested in what is happening close to the airfoil, the output time can be reduced by a factor two to three which reduces the simulation time by approximately the same factor. Eventually, $C_{p}$ values are obtained at the wall using the inverse distance weighting (order two) of values available on the nearest and second nearest fluid cells.
@@ 6.5 Partial conclusions
This validation has further confirmed that our approach, which combines (1) a high-order formulation of conservative CC grid refinement, (2) stair-cased bounce-back techniques, (3) numerical equilibria, and (4) kinetic-based shock sensors is both accurate and robust for fully compressible flow simulations. The generality of our framework also facilitated the implementation of the grid refinement algorithm for several high-order lattices (D2Q13, D2Q21, and D3Q37). This significantly broadens the operational range of the initial low-speed solver so that transonic, supersonic, and potentially higher-speed regimes can be simulated.
@@ 7.1 Compilation flags, optimization levels and performance measurements
The -O1 optimization level was chosen after comparison with higher levels (-O2, -O3, -Ofast, and -O4). While these higher levels may improve performance in some cases, they often led to performance degradation. Notably, we found out that the more aggressive optimizations (-O3 and above) could cause up to $20\%$ performance losses for specific versions of the compiler. For the 24.1 version, -O1 was chosen as a good tradeoff between code optimization and performance. Given the frequent updates to the nvc++ compiler (6 to 7 releases per year), it is recommended to regularly perform performance analyses to prevent regressions when upgrading to newer versions.
@@
For all benchmarks, performance is averaged over three consecutive full simulations through
@@ 7.2 Performance on matrix-like grids
Here, $n_{LBM}$ represents the number of variables stored in memory for a given LB scheme. Specifically, $n_{LBM}=Q$ and $2Q+n_{\lambda}$ for weakly and fully compressible models, respectively. Each variable requires two memory accesses per time step –one for reading it and one for writing it back to memory. In this work, four different lattices are used ($Q\in\{9,13,21,37\}$), and the number of Lagrange multipliers is $n_{\lambda}=8$. $s_{\mathrm{data}}=4$ or $8$ bytes depending on the arithmetic precision used for the simulations, and two flags of $4$ bytes each ($s_{\mathrm{index}}$ and $s_{\mathrm{type}}$) are used for the cell index and its type (fluid, solid, etc). In practice, one can expect weakly compressible LBMs to reach 60 to 80% of the peak performance on GPUs using modern C++ [11, 12], while fully compressible LBMs used in this work perform at about 40% for the most powerful GPUs [51]. Achieving this level of performance requires extensive code optimization, including techniques such as loop unrolling and precomputing redundant terms. Additionally, it necessitates sufficiently high memory utilization, typically ranging from a few percent to up to 20% of the total memory for server-based and gaming-class GPUs, respectively.
Fig 22: Performance analysis of the lid-driven cavity simulation using a matrix-like SoA implementation of the D2Q9-LBM with various collision models and $(\mathrm{Re}$,$\mathrm{Ma})=(100,=0.2)$. For the lowest cavity resolution, $L=2^{7}=128$ cells per direction, the performance is strongly impacted by thread over-synchronization. As $L$ increases, all models progressively reach 80% of the peak performance of the A100 SXM4 (80GB) that is computed using the performance model (14). The Reynolds number is increased to 10’000 when $L\geq 2^{10}=1024$ to reduce the value of $\tau$ and avoid stability issues.
To support our arguments, we included in our framework an additional implementation of the lid-driven cavity benchmark using matrix-like data structures with an SoA memory layout and “two-population” storage formalism (see Section 4.1). Performance data obtained with the D2Q9 lattice and several collision models (BGK, TRT and RR) are reported on Figure 22. As previously discussed, the performance barely depends on the collision model, and it is an increasing function of the cavity length resolution $L$. More precisely, it reaches a plateau (about $80$% of the peak performance), when $L$ is greater than $40962^{12}=$ 4096 cells per direction, i.e., a few percent of the memory size of the A100 SXM4 (80GB). This benchmark also confirms that there is no need to rely on low-level coding languages, like CUDA, to achieve near-peak performance on GPUs with matrix-like LBMs.
7.3 Performance on tree-like structures with uniform grids
@@
Table 1: Performance ratios $P_{\mathrm{ratio}}$ (16) computed for various LBMs and arithmetic precisions. sp and dp stands for single precision and double precision, respectively.
@@
Building on this theoretical performance analysis, we further evaluate the performance of our framework using the lid-driven cavity benchmark with a uniform grid stored in a tree-like structure. As in Section 7.2, the simulations are conducted with the D2Q9-LBM and various collision models (BGK, TRT, RR). Figure 23 presents results for both single-precision and double-precision arithmetic. Consistent with matrix-like grids, (1) performance remains rather independent of the collision model, and (2) it increases with memory usage, reaching a plateau for $L\geq 2^{12}$. However, the peak performance reaches only 60% of the theoretical maximum. This limitation and the difference between collision models are attributed to the partial optimization of our framework, which prioritizes the validation of various LBMs rather than optimization for a specific model. Nevertheless, the performance of the D2Q9-RR ranks among the best reported in the literature, notably those obtained with highly optimized approaches based on trees of blocks and GPU-specific coding languages like CUDA and HIP [10]. Similar trends are observed for other weakly compressible LBMs employed in this study, such as the D2Q13-RR and D2Q21-RR, which achieve approximately 45% of the peak performance (see B.7).
Fig 23: Performance analysis of the lid-driven cavity simulation using a tree-like SoA implementation of the D2Q9-LBM with various collision models and $(\mathrm{Re}$,$\mathrm{Ma})=(100,=0.2)$. For the lowest cavity resolution, $L=2^{7}=128$ cells per direction, the performance is strongly impacted by thread over-synchronization. As $L$ increases, all models progressively reach 60% of the peak performance of the A100 SXM4 (80GB) that is computed using the performance model (15). The Reynolds number is increased to 10’000 when $L\geq 2^{10}=1024$ to reduce the value of $\tau$ and avoid stability issues.
7.4 Performance on tree-like structures with non-uniform grids
@@
$P_{\mathrm{peak}}^{\mathrm{tree},*}=\dfrac{1}{\Delta t_{c}}\sum_{t=0}^{\Delta t_{c}}\left[\sum_{i=0}^{i_{max}}P_{\mathrm{peak}}^{\mathrm{tree}}(i)\dfrac{\delta_{2^{i}}(t)}{||\delta_{2^{i}}(t)||}\right],$ (19)
with $\delta_{2^{i}}(t)$ being the Kronecker function that returns one when level $i$ requires an update (every $2^{i}$ time iterations), and zero otherwise. $||\delta_{2^{i}}(t)||$ is defined as the number of times $\delta_{2^{i}}(t)=1$ when looping over all refinement levels during $\Delta t_{c}$.
@@ 7.5 Concrete impact of the grid hierarchy on performance: Aeolian noise simulations
In Table 2, two key metrics are the total number of cells in the grid, and the equivalent fine cells. The former quantifies the memory footprint of the simulation, whereas the latter quantifies the computational savings achieved by the asynchronous grid refinement strategy. Specifically, it is derived by summing all cells and weighting them by their update frequency ($1/2^{i}$). By doing so, one can measure how much of the computational load is concentrated in finer grid areas, while coarser areas are updated less frequently. Additionally, the table includes a configuration with a uniform grid consisting solely of fine cells. This grid requires $2^{\text{depth}}$ cells per direction, corresponding to a total cell count of 67’108’864. This configuration is mainly used to quantify the differences in computational load and memory requirements between grids with and without refinement. Additional discussions about the parametric studies can be found in B.8.
@@
    Configurations with the highest update frequencies do not correspond to those with the shortest simulation times. Instead, near-peak performance is observed for the largest grids (#5, #6, #11, and #12), which, conversely, result in the longest simulation times (#1, #7, #2, and #3).
The third observation is particularly important as it highlights the tradeoff required when performing simulations on non-uniform grids using GPUs. To end up with the shortest simulations, it is essential to minimize the number of cells in the finest refinement levels that are updated the most frequently. Based on that, configurations #4 and #5 lead to optimal acoustic propagation in the far-field, since they have the best performance-to-simulation-time ratio while keeping a high mesh resolution far away from the cylinder. Conversely, configurations #8 and #9 lead to optimal near-field aerodynamics with good far-field accuracy, since these grids keep a high mesh resolution close to the cylinder while refining the mesh far away from the cylinder. Yet, this comes at the cost of larger memory footprints compared to #4 and #5. Similar trends are observed for grids with different depths (see B.8).
In conclusion, adopting best practices for the grid generation is essential to balancing accuracy, performance, and simulation time effectively. These decisions are inherently case-dependent, as the priority between near-field aerodynamics and far-field acoustics dictates the appropriate grid configuration. The performance model, and the insights from this case study, provide a solid foundation for understanding all the aspects of such problems and for deriving best practices in a semi-automated manner.
@@ 7.6 Preliminary performance results on fully compressible models
    Due to the compute-bound nature of our compressible LBMs, increasing the lattice size leads to additional performance loss compared to memory-bound LBMs. For instance, both matrix-like and tree-like implementations of the D2Q13 and D2Q21 RR-LBM exhibit a performance ratio of $P_{Q21}/P_{Q13}\approx 13/21\approx 62\%$ (see Figures 41 and 42). Hence, moving from 13 to 21 velocities leads to a performance loss is 38%. However, in the fully compressible case, the performance loss reaches about 50%, indicating an additional 10% performance loss. Interestingly, this 10% loss remains consistent across all lattices and grid configurations considered in this work.
@@ 8.2 Testcase presentation, motivation, and results
Fig 27: Illustration of AMR for the supersonic flow past a NACA airfoil with D2Q21-LBM. Time evolution of the grid (top), and Mach field (bottom) highlight how adaptive refinement patches track the primary bow shock while seamlessly integrating with fixed patches.
The convected vortex benchmark evaluates the accuracy of AMR when refinement patches follow a translating and rotating vortex. A circular refinement region encloses the vortex, with two perpendicular rectangular patches forming a ‘propeller-like’ structure that moves and rotates synchronously with the vortex. As the mesh motion continuously refines and coarsens cells upstream and downstream of the vortex, this test helps quantifying whether the basic coalescence and uniform explosion steps introduce round-off errors that could significantly distort vortex evolution. In this work, the vortex is initialized in a periodic domain of size $L$, using a barotropic formulation (see Eq. (29) in Ref. [205]) with a radius $R=L/40$. The Mach number is 0.2, and the Reynolds number is chosen very high ($\mathrm{Re}=10^{9}$) so that possible errors originating from coarsening and refining cannot be attenuated by viscous effects. Velocity fields are reported on the top panels of Figure 26 for different dimensionless times. They show that the motion of the vortex is not impacted by the dynamic remeshing happening in front and behind it, as it crosses the full domain five times.
@@
9 General conclusions and perspectives
Despite more than 30 years of research and development, designing lattice Boltzmann methods (LBMs) that run accurately, robustly, and efficiently on non-uniform grids while allowing seamless GPU acceleration remains a challenge. This difficulty stems from the fact that both academic and applied studies often focus on these aspects in an isolated manner instead of addressing them all at once. This motivated us to propose a strategy that couples simple, yet effective ideas to achieve (1) seamless hardware acceleration, (2) accurate and robust simulations across a wide range of Reynolds and Mach numbers, and (3) easy handling of complex geometries encountered in industrial applications.

### v3

@@ 6.3 Viscous flow past a NACA0012 airfoil at transonic speeds
To tackle this configuration, we use a simulation domain of size $\left[-15C,15C\right]^{2}$ -with $C$ the airfoil chord- that is bounded by three equilibrium Dirichlet boundary conditions (left, top, and bottom) and one Neumann condition (right). As a best practice from previous works [67, 51], we start with a uniform flow at a freestream Mach number lower than the targeted one. The freestream Mach number imposed by the Dirichlet boundary conditions is then rapidly increased (over one convective time) to its nominal value. This notably allows to avoid convergence issues that were encountered with the Newton-Raphson algorithm at the beginning of the simulation for the D2Q13-LBM. When it comes to the mesh generation, four rectangular and three NACA-shaped refinement patches surround the airfoil. The finest refinement patch aims at (1) better capturing the boundary layer development at the leading edge of the airfoil, and (2) reducing pressure oscillations at the wall that are partly induced by the staircase nature of the bounce-back approach. Hereafter, simulation data are outputted at $t=5t_{c}^{\mathrm{domain}}$, so that the transient flow is completely evacuated from the simulation domain, and with the characteristic time being $t_{c}^{\mathrm{domain}}=30C/(\mathrm{Ma}\sqrt{\gamma_{r}T_{0}})$. If one is only interested in what is happening close to the airfoil, the output time can be reduced by a factor two to three which reduces the simulation time by approximately the same factor. Eventually, $C_{p}$ values are obtained at the wall using the inverse distance weighting (order two) of values available on the nearest and second nearest fluid cells.
@@ 6.5 Partial conclusions
This validation has further confirmed that our approach, which combines (1) a high-order formulation of conservative CC grid refinement, (2) stair-cased bounce-back techniques, (3) numerical equilibria, and (4) kinetic-based shock sensors is both accurate and robust for fully compressible flow simulations. The generality of our framework also facilitated the implementation of the grid refinement algorithm for several high-order lattices (D2Q13, D2Q21, and D2Q37). This significantly broadens the operational range of the initial low-speed solver so that transonic, supersonic, and potentially higher-speed regimes can be simulated.
@@ 7.1 Compilation flags, optimization levels and performance measurements
The -O1 optimization level was chosen after comparison with higher levels (-O2, -O3, -Ofast, and -O4). While these higher levels may improve performance in some cases, they often led to performance degradation. Notably, we found out that the more aggressive optimizations (-O3 and above) could cause up to $20\%$ performance losses for specific versions of the compiler. For the 24.1 version, -O1 was chosen as a good tradeoff between code optimization and performance. Given the frequent updates to the nvc++ compiler (6 to 7 releases per year), it is recommended to regularly run performance analyses to prevent regressions when upgrading to newer versions.
@@
For all benchmarks, performance is averaged over three consecutive full simulations, and it is computed as
@@ 7.2 Performance on matrix-like grids
Here, $n_{\mathrm{LBM}}$ represents the number of variables stored in memory for a given LB scheme. Specifically, $n_{\mathrm{LBM}}=Q$ and $2Q+n_{\lambda}$ for weakly and fully compressible models, respectively. Each variable requires two memory accesses per time step –one for reading it and one for writing it back to memory. In this work, four different lattices are used ($Q\in\{9,13,21,37\}$), and the number of Lagrange multipliers is $n_{\lambda}=8$. $s_{\mathrm{data}}=4$ or $8$ bytes depending on the arithmetic precision used for the simulations, and two flags of $4$ bytes each ($s_{\mathrm{index}}$ and $s_{\mathrm{type}}$) are used for the cell index and its type (fluid, solid, etc). In practice, one can expect weakly compressible LBMs to reach 60 to 80% of the peak performance on GPUs using modern C++ [11, 208], while fully compressible LBMs used in this work perform at about 40% for the most powerful GPUs [51]. Achieving this level of performance requires extensive code optimization, including techniques such as loop unrolling and precomputing redundant terms. Additionally, it necessitates sufficiently high memory utilization, typically ranging from a few percent to up to 20% of the total memory for server-based and gaming-class GPUs, respectively.
Fig 22: Performance analysis of the lid-driven cavity simulation using a matrix-like SoA implementation of the D2Q9-LBM with various collision models and $(\mathrm{Re}$$\mathrm{Ma})=(100,=0.2)$. For the lowest cavity resolution, $L=2^{7}=128$ cells per direction, the performance is strongly impacted by thread over-synchronization. As $L$ increases, all models progressively reach 80% of the peak performance of the A100 SXM4 (80GB) that is computed using the performance model (14). The Reynolds number is increased to 10’000 when $L\geq 2^{10}=1024$ to reduce the value of $\tau$ and avoid stability issues.
To support our arguments, we included in our framework an additional implementation of the lid-driven cavity benchmark using matrix-like data structures with an SoA memory layout and “two-population” storage formalism (see Section 4.1). Performance data obtained with the D2Q9 lattice and several collision models (BGK, TRT and RR) are reported on Figure 22. As previously discussed, the performance barely depends on the collision model, and it is an increasing function of the cavity length resolution $L$. More precisely, it reaches a plateau (about 8080% of the peak performance), when $L$ is greater than $2^{12}=4096$ cells per direction, i.e., a few percent of the memory size of the A100 SXM4 (80GB). This benchmark also confirms that there is no need to rely on low-level coding languages, like CUDA, to achieve near-peak performance on GPUs with matrix-like LBMs.
7.3 Performance on tree-like structures with uniform grids: Quantifying the impact of the neighbor-list
@@
Table 1: Theoretical performance ratios $P_{\mathrm{ratio}}$ (16) computed for various LBMs and arithmetic precisions. sp and dp stands for single precision and double precision, respectively.
@@
Building on this theoretical performance analysis, we further evaluate the performance of our framework using the lid-driven cavity benchmark with a uniform grid stored in a tree-like structure. As in Section 7.2, the simulations are conducted with the D2Q9-LBM and various collision models (BGK, TRT, RR). Figure 23 presents results for both single-precision and double-precision arithmetic.
Fig 23: Performance analysis of the lid-driven cavity simulation using a tree-like SoA implementation of the D2Q9-LBM with various collision models and $(\mathrm{Re}$$\mathrm{Ma})=(100,=0.2)$. For the lowest cavity resolution, $L=2^{7}=128$ cells per direction, the performance is strongly impacted by thread over-synchronization. As $L$ increases, all models progressively reach 60% of the peak performance of the A100 SXM4 (80GB) that is computed using the performance model (15). The Reynolds number is increased to 10’000 when $L\geq 2^{10}=1024$ to reduce the value of $\tau$ and avoid stability issues.
Consistent with matrix-like grids, (1) performance remains rather independent of the collision model, and (2) it increases with memory usage, reaching a plateau for $L\geq 2^{12}$. However, the peak performance reaches only 60% of the theoretical maximum. This value and the difference between collision models are attributed to the partial optimization of our framework, which prioritizes the validation of various LBMs rather than optimization for a specific model. Nevertheless, the performance of the D2Q9-RR ranks among the best reported in the literature, notably those obtained with highly optimized approaches based on trees of blocks and GPU-specific coding languages like CUDA and HIP [10]. In the latter case, the performance is roughly halved, compared to simpler matrix-based formulations. This is in agreement with the performance ratios measured with our framework (see Table 2), even if they are approximately 20% lower than our theoretical estimates (Table 1). Nevertheless, they remain sufficiently close to lend confidence to the validity of our performance model.
Single Precision (GLUPS) Double Precision (GLUPS)
Collision Matrix Tree $P_{\mathrm{ratio}}$ Matrix Tree $P_{\mathrm{ratio}}$
BGK 20.503 10.739 0.52 10.598 5.824 0.55
TRT 20.390 10.688 0.52 10.178 5.847 0.57
$R$ 20.584 10.541 0.51 10.795 7.041 0.65
Table 2: Measured performance ratios $P_{\mathrm{ratio}}$ for D2Q9-LBMs. Data is gathered from lid-driven cavity simulations using $L=2^{13}$.
Similar trends in peak performance and impact of neighbor-list are observed for other weakly compressible LBMs used in this study, such as the D2Q13-RR and D2Q21-RR. For these high-order LBMs, the performance results remain promising and suggest clear opportunities for further optimization (see B.7).
7.4 Performance on tree-like structures with non-uniform grids: Theoretical impact of asynchronous time-stepping
@@
[…]
with $\delta_{2^{i}}(t)$ being the Kronecker function that returns one when level $i$ requires an update (every $2^{i}$ time iterations), and zero otherwise. $\delta_{2^{i}}(t)||$ is defined as the number of times $\delta_{2^{i}}(t)=1$ when looping over all refinement levels during $\Delta t_{c}$.
@@ 7.5 Concrete impact of the grid hierarchy on performance: Aeolian noise simulations
In Table 3, two key metrics are the total number of cells in the grid, and the equivalent fine cells. The former quantifies the memory footprint of the simulation, whereas the latter quantifies the computational savings achieved by the asynchronous grid refinement strategy. Specifically, it is derived by summing all cells and weighting them by their update frequency ($1/2^{i}$). By doing so, one can measure how much of the computational load is concentrated in finer grid areas, while coarser areas are updated less frequently. Additionally, the table includes a configuration with a uniform grid consisting solely of fine cells. This grid requires $2^{\text{depth}}$ cells per direction, corresponding to a total cell count of 67’108’864. This configuration is mainly used to quantify the differences in computational load and memory requirements between grids with and without refinement. Eventually, the ratio between the total number of cells and the equivalent fine cells can be interpreted as the theoretical speedup provided by asynchronous time stepping, under the ideal assumption that performance is independent of memory usage. Specifically, under iso-LUPS conditions, configuration #1 achieves a speedup of approximately $23114/1338\approx 17$, thanks to asynchronous time stepping. Additional discussions about the parametric studies can be found in B.8.
@@
    Configurations with the highest update frequencies do not correspond to those with the shortest simulation times. Instead, near-peak performance is observed for the largest grids (#6, #11, and #12), which, conversely, result in the longest simulation times. Opposite conclusions are reported for the smallest grids (#1, #2, and #7).
The third observation is particularly important as it highlights the tradeoff required when performing simulations on non-uniform grids using GPUs. To minimize simulation time, two key factors must be considered: (1) reducing the number of cells at the finest refinement levels, as these are updated most frequently, and (2) maintaining a sufficiently high global update rate (LUPS). Based on these principles, configurations #4 and #5 provide optimal far-field acoustic propagation. They offer the best performance-to-simulation-time trade-off, while also preserving high mesh resolution in the far field, which is critical for accurately capturing acoustic wave propagation. Conversely, configurations #9 and #10 lead to optimal near-field aerodynamics with good far-field accuracy, since these grids keep a high mesh resolution close to the cylinder while coarsening the mesh far away from the cylinder. Yet, this comes at the cost of larger memory footprints compared to #4 and #5. Similar trends are observed for grids with higher mesh resolutions, except for the normalized performance-to-simulation-time ratio, which consistently improves with increasing resolution due to the higher workload across all refinement levels (see Appendix B.8).
In conclusion, adopting best practices for the grid generation is essential to balancing accuracy, performance, and simulation time effectively. These decisions are inherently case-dependent, as the priority between minimal simulation time, memory footprint, accurate near-field aerodynamics, or accurate far-field acoustics dictates the appropriate grid configuration. The performance model, and the insights from this case study, provide a solid foundation for understanding all the aspects of such problems and for deriving best practices in a semi-automated manner.
@@ 7.6 Preliminary performance results on fully compressible models
    Due to the compute-bound nature of our compressible LBMs, increasing the lattice size leads to additional performance loss compared to memory-bound LBMs. For instance, both matrix-like and tree-like implementations of the D2Q13 and D2Q21 RR-LBM exhibit a performance ratio of $P_{Q21}/P_{Q13}\approx 13/21\approx 62\%$ (see Figures 42 and 43). Hence, moving from 13 to 21 velocities leads to a performance loss is 38%. However, in the fully compressible case, the performance loss reaches about 50%, indicating an additional 10% performance loss. Interestingly, this 10% loss remains consistent across all lattices and grid configurations considered in this work.
@@ 8.2 Testcase presentation, motivation, and results
Fig 27: Illustration of AMR for the supersonic flow past a NACA airfoil with D2Q21-LBM. Time evolution of the grid (top), and Mach field (bottom) highlight how adaptive refinement patches track the primary bow shock while seamlessly merging with fixed patches.
The convected vortex benchmark evaluates the accuracy of AMR when refinement patches follow a translating and rotating vortex. A circular refinement region encloses the vortex, with two perpendicular rectangular patches forming a ‘propeller-like’ structure that moves and rotates synchronously with the vortex. As the mesh motion continuously refines and coarsens cells upstream and downstream of the vortex, this test helps quantifying whether the basic coalescence and uniform explosion steps introduce round-off errors that could significantly distort vortex evolution. It also provides insight into whether this strategy could be used in the future to simulate scenarios involving moving/rotating body forces, such as those in actuator models commonly used to replicate the effects of drone or wind turbine blades without explicitly modeling them (see, e.g., Ref. [205]). In this work, the vortex is initialized in a periodic domain of size $L$, using a barotropic formulation (see Eq. (29) in Ref. [206]) with a radius $R=L/40$. The Mach number is 0.2, and the Reynolds number is chosen very high ($\mathrm{Re}=10^{9}$) so that possible errors originating from coarsening and refining cannot be attenuated by viscous effects. Velocity fields are reported on the top panels of Figure 26 for different dimensionless times. They show that the motion of the vortex is not impacted by the dynamic remeshing happening in front and behind it, as it crosses the full domain five times.
@@
9 Synthesis of findings
In this work, we propose a strategy that couples simple, yet effective ideas to achieve (1) seamless hardware acceleration, (2) accurate and robust simulations across a wide range of Reynolds and Mach numbers, and (3) easy handling of complex geometries encountered in industrial applications.